OpenStudy (anonymous):

Find all orders pairs (x, y) that satisfy the system:

OpenStudy (anonymous):

\[x ^{3}-x ^{2}y+5xy ^{2}-y ^{3}=124\] \[x ^{3}-5x ^{2}y+xy ^{2}-y ^{3}=4\]

OpenStudy (anonymous):

ok finally got it :)

OpenStudy (anonymous):

x3−x2y+5xy2−y3=124 R1 x3−5x2y+xy2−y3=4 R2 R1-R2 => -x2y+5xy2+5x2y-xy2 = 120 (-x2y - xy2) + (5xy2 + 5x2y) =120 -xy(x+y) + 5xy(x+y) =120 4xy(x+y)=120 (*) no lets go back to R1 x3 - x2y + 5xy2 - y3 =124 x3 - 3x2y +2x2y + 3xy2 + 2xy2 - y3 = 124 (x3 - 3x2y + 3xy2 -y3) + ( 2x2y + 2xy2) =124 (x-y)3 + 2xy(x+y) =124 (**) replace (*) into (**) (x-y)3 + 60 = 124 (x-y)3=64 x-y = 4 x=4+y repalce this back into (*) 4(4+y)(4+y+y)=120 then solve for y

OpenStudy (anonymous):

Doesn't R1-R2 => 4x2y+4xy2=120

OpenStudy (anonymous):

oh yes, my bad, u're right

OpenStudy (anonymous):

I did the whole thing out and got \[x=\sqrt{30\div(y+y ^{2})}\]

OpenStudy (anonymous):

How does that help when I plug it back into the first equation?

OpenStudy (anonymous):

4(4+y)(4+y+y)=120 8(y2+6y+8)=120 y2+6y+8=15 solve for y then u get y=-2 and y=-4